While doing their homework together the other night, Jack and Jill came upon the following question: Find the equation of the line that is parallel to the line
3x + 7y = 10 and passes through the point (4, 8). Jack said that it was impossible to solve this. Jill told Jack that he was just lazy, and of course you can get an answer. Jill then proceeded to show Jack how. Was Jill right? Could she solve this? If so, show the steps that Jill used to prove Jack wrong. If you think Jack is correct, explain why.

Answer by jim_thompson5910(35256)

It's possible to solve this.

Anything parallel to 3x + 7y = 10 is of the form 3x + 7y = C where C is some constant.

Now plug in (x,y) = (4,8)

3x + 7y = C

3(4) + 7(8) = C

12 + 56 = C

68 = C

C = 68

So the equation of the parallel line is 3x + 7y = 68 (note: this is in standard form)